Three years ago a father was thrice as old as his son in 5 years time the father will be twice as his son what will be the sum of their age in four years
Answer by ankor@dixie-net.com(22740)

let f = father's present age, let s = son's
:
write an equation for each statement, simplify
:
Three years ago a father was thrice as old as his son
f - 3 = 3(s-3)
f - 3 = 3s - 9
f = 3s - 9 + 3
f = 3s - 6
in 5 years time the father will be twice as his son
f + 5 = 2(s+5)
f + 5 = 2s + 10
f = 2s + 10 - 5
f = 2s + 5
replace f with 3s - 6, from the 1st statement
3s - 6 = 2s + 5
3s - 2s = 5 + 6
s = 11 yrs is the son's age
find the fathers age
f = 3(11) - 6
f = 33 - 6
f = 27 yrs is the father's age
:
what will be the sum of their age in four years
27+4 + 11+4 = 46